Factor the given polynomial completely. If the polynomial cannot be​ factored, say that it is prime.


x2 - x -56


Is this prime, if not what is it equal to? please show work



1
Expert's answer
2021-09-28T01:27:39-0400

x2 - x - 56 = 0


x2 - 8x + 7x -56 = 0

x(x-8) + 7(x-8) = 0

(x-8)(x+7) = 0


x-8 = 0 and x+7 = 0


x= 8 and x = -7